Write the linear equation in the form y=mxtb. identify the slope and y intercept: 2xt3y-10 my teacher's answer was :
y=mxtb
3y/3= -2xt10/3
y= -2x/3 + 10/3
m= -2/3
y= 10/3
is he correct?

Answer by Greenfinch(383)

It will work better as y = mx + b
Equation is 3y + 2x = 10
divide by 3, y =- 2/3 x + 10/3
Gradient is co-efficient of x or -2/3
y-intercept is 10/3